How much heat is required to heat 50.0 g of water from 25.0 ̊C to 75.0 ̊C?
1
Expert's answer
2020-12-18T08:46:49-0500

q = mC∆T

∆T = (75-25) = 50°C

m = 50g

C = 4.18 J/g/°C


q = 50× 4.18× 50

q = 10450 J which is equal to 10.45KJ